1) Given the following conditional value table, determine the appropriate decision under uncertainty using :

a) Maximax

b) Maximin

c) Equally likely

                                                                                               States of Nature

Alternatives

Very Favorable Market

Average Market

Unfavorable Market

Large Plant

   $275,000

   $100,000

   -$150,000

Small Plant

$200,000

   $60,000

-$10,000

Overtime

   $100,000

    $40,000

   -$1000

Do Nothing

   $0

   $0

   $0
Answer #1

Answer:

Maximax:

  • Under maximax we choose alternative having the highest maximum payoff
  • MAx payoffs:
    • Large plant = 275,000
    • Small plant = 200000
    • Overtime= 100000
    • Do nothing = 0
  • Hence we will choose large plant in Mximax strategy with the max payoff of $275,000

Maximin:

  • Under maximin we choose alternative having the highest minimum payoff
  • Min payoffs:
    • Large plant = -150,000
    • Small plant = -10000
    • Overtime= -1000
    • Do nothing = 0
  • Hence we will choose do nothing in Minimax strategy with min payoff of 0.

Equally, like:

  • Under equally likely we choose alternative having the highest average payoff.
  • average payoffs:
    • Large plant = (275,000+100000-150000)/3 = $75,000
    • Small plant = (200000+60000-10000)/3 = $50,000
    • Overtime= (100000+40000-1000)/3 =$46333.3
    • Do nothing = 0
  • Hence we will choose large plant in Equally likely with average payoff of $75,000